Solve the PDE dz/dx+dz/dy = z^2


1
Expert's answer
2021-04-29T17:57:53-0400

Solve the PDE dz/dx+dz/dy = z^2

Solution:

A parametrization invariant is:

"dx=dy=\frac{dz}{z^2}"

From "dx = du" ,

"x-y=C_1" which is a first characteristic equation.

From "dy=\frac{dz}{z^2}" ,

"y=-\frac{1}{z}+C_2" is the second characteristic equation.

The general solution of the PDE expressed on the form of an implicit equation is :

"\Phi((x-y),(y+\frac1z))=0"

where "\Phi" is any differentiable function of two variables.

Answer: "\Phi((x-y),(y+\frac1z))=0" .